OpenStudy (gabylovesyou):

Which set of data shows 40% as the experimental probability of a particular traffic light being green on your way to school in the morning? A. spinning 20 vowels out of 40 spins of a spinner with 6 equally divided sections labeled with 3 vowels and 3 consonants B. 2 tails in 20 coin flips, with tails representing a green light C. 2 odd-numbers in 8 rolls of a standard die, with odd numbers representing a green light D. 10 of 25 selections at random from a card deck being a black card, with black representing a green light

OpenStudy (gabylovesyou):

@asnaseer @alexwee123 @Abbie23 @alexeis_nicole @brinethery @breebree8 Help? thanks

OpenStudy (anonymous):

kay so first to get the answer you must divide the numbers given to you. A. 20/40 = 0.5 B. 2/20=0.1 C. 2/8=0.25 D. 10/25= 0.40 so.. now given those numbers which one is 40%

OpenStudy (gabylovesyou):

D

OpenStudy (anonymous):

yep that's right.

OpenStudy (gabylovesyou):

OMG thank you soooooo much!!!!!!!!!!!!!

OpenStudy (anonymous):

np. :)